# Vacuum amplitudes and time-like causal unitary in the loop-tree duality

### v1

@@ Abstract
We present the first proof-of-concept application to decay processes at higher perturbative orders of LTD causal unitary, a novel methodology that exploits the causal properties of vacuum amplitudes in the loop-tree duality (LTD) and is directly well-defined in the four physical dimensions of the space-time. The generation of loop- and tree-level contributions to the differential decay rates from a kernel multiloop vacuum amplitude is shown in detail, and explicit expressions are presented for selected processes that are suitable for a lightweight understanding of the method. Specifically, we provide a clear physical interpretation of the local cancellation of soft, collinear and unitary threshold singularities, and of the local renormalisation of ultraviolet singularities. The presentation is illustrated with numerical results that showcase the advantages of the method.
I Introduction
In a recent paper [1], we have proposed a novel representation of differential observables at high-energy colliders that exploits the manifestly causal properties of the loop-tree duality (LTD) at higher perturbative orders and its connections with directed acyclic graph (DAG) configurations in graph theory [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 11, 14, 15, 16, 17]. This approach, dubbed LTD causal unitary, generalises the method of four-dimensional unsubtraction (FDU) [18, 19, 20, 21, 22, 23, 24, 25] in which, unlike subtraction methods [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67], infrared (IR) and ultraviolet (UV) singularities are locally cancelled directly between loop and tree contributions at the integrand level, while introducing relevant improvements and new features that facilitate an efficient implementation.
The central and novel ingredient of LTD causal unitary is a multiloop vacuum amplitude in the LTD representation that depends on $\Lambda$ loop momenta, $\{\ell_{j}\}_{j=1,\ldots,\Lambda}$. In LTD causal unitary, the differential contribution to the decay rate of a particle of mass $m_{a}$ from the $k^{\rm{th}}$-order in perturbation theory is given by
$d\Gamma^{(k)}_{a}=\frac{d\Lambda}{2m_{a}}\,\sum_{(i_{1}\cdots i_{n}a)\in\Sigma}{\cal A}_{\rm D}^{(\Lambda,{\rm R})}(i_{1}\cdots i_{n}a)\,\widetilde{\Delta}_{i_{1}\cdots i_{n}\bar{a}}~{},$
@@ I Introduction
where $d\Gamma^{{\rm N}^{k}{\rm LO}}_{a}$ denotes the differential decay rate up to (next-to)${}^{k}$-leading order, and the integration measure
@@
The final states with $n$ particles are then generated from residues, called phase-space residues, of a vacuum amplitude ${\cal A}_{\rm D}^{(\Lambda)}$ on the on-shell energies of the internal propagators, $q_{i_{s},0}^{(+)}=\sqrt{\mathbf{q}_{i_{s}}^{2}+m_{i_{s}}^{2}-\imath 0}$, where $\imath 0$ is the original complex prescription of a Feynman propagator, $\mathbf{q}_{i_{s}}$ the spacial components of the four-momentum $q_{i_{s}}$, and $m_{i_{s}}$ its mass,
${\cal A}_{\rm D}^{(\Lambda,{\rm R})}(i_{1}\cdots i_{n}a)$ $=$ ${\cal A}_{\rm D}^{(\Lambda)}(i_{1}\cdots i_{n}a)$ (4)
-- ${\cal A}_{{\rm UV}}^{(\Lambda)}(i_{1}\cdots i_{n}a)~{},$
@@
The LTD causal unitary representation in Eq. (1) involves the sum over all possible phase-space residues of the vacuum amplitude. It is precisely the unitary sum over all of them that ensures that most of the unique properties of the vacuum amplitude are preserved. Specifically, the vacuum amplitude in LTD is a function of the on-shell energies and is obtained by replacing the Feynman propagators by causal propagators of the form
@@
Since the real part of the on-shell energies is positive by definition, the causal propagators in Eq. (9) cannot become singular. If all the on-shell energies vanish simultaneously for massless particles, the potential soft singularity is screened by the integration measure which also vanishes in this limit. As a consequence, the vacuum amplitude cannot exhibit soft or collinear singularities, and remarkably unitary threshold singularities are also absent. The absence of singularities in the vacuum amplitude, apart from UV singularities that are accounted for by the local UV counterterm, ensures that the unitary sum over the phase-space residues is also free of soft, collinear and unitary threshold singularities [1]. Therefore, the master expression for the decay rate in Eq. (1) is well defined in the four physical dimensions of the space-time, since all potential singularities locally cancel out between different phase-space residues.
@@
The outline of the paper is as follows. In Section II, we present explicit expressions for the phase-space residues that contribute to the decay processes $H\to q\bar{q}(g)$ and $\gamma^{*}\to q\bar{q}(g)$ at LO and NLO. In Section III, we construct the corresponding UV counterterms, which implement a local UV renormalisation, and discuss in detail the wave-function renormalisation and the mass renormalisation scheme. In Section IV, we consider a simplified model that has all the main properties of a NNLO calculation and is therefore suitable for presenting LTD causal unitary at NNLO in a lightweight form. In Section V, we present a numerical implementation of LTD causal unitary at NLO and NNLO based on the phase-space residues presented in Sections II to IV, and illustrate the local cancellation of unitary threshold, double- and triple-collinear/quasicollinear singularities. Finally, in Section VI we draw our conclusions and future prospects.
II Vacuum amplitudes and decay rates at NLO
@@
$q_{1}=\ell_{1}+\ell_{2}~{},\quad q_{2}=\ell_{1}+\ell_{3}~{},\quad q_{3}=\ell_{1}~{},$
$q_{4}=\ell_{2}~{},\quad q_{5}=\ell_{2}-\ell_{3}~{},\quad q_{6}=\ell_{3}~{}.$ (10)
We use the shorthand notation $\ell_{ij}=\ell_{i}+\ell_{j}$ and $\ell_{i\bar{j}}=\ell_{i}-\ell_{j}$. Nonetheless, it is worth noting that one of the advantages of the LTD representation is that it is independent of the specific labelling of the internal momenta. The momentum $q_{1}$ represents a gluon and is massless, while $q_{2}$ through $q_{5}$ are quarks or antiquarks of mass $m$. The momentum $q_{6}$ is either a Higgs boson or an off-shell photon. To simplify the presentation, we define $x_{i_{1}\cdots i_{n}}=\prod_{s=1}^{n}2q_{i_{s},0}^{(+)}$.
@@
For convenience, the average factor over initial-state polarizations is included in ${\cal A}_{\rm D}^{(2,f)}$. In this expression, we have anticipated that $\overline{|{\cal M}_{f\to q\bar{q}}^{(0)}|^{2}}$ is the matrix element squared obtained with the standard approach of squaring the tree-level amplitude for $f\to q\bar{q}$ and averaging over the initial-state polarizations, where the interaction couplings and final-state colour factors are collected in the coefficient $g_{f}^{(0)}$,
@@
$\Gamma^{\rm LO}_{H\to q\bar{q}}=g_{H}^{(0)}\frac{\beta^{3}m_{H}}{8\pi}~{},$
$\Gamma^{\rm LO}_{\gamma^{*}\to q\bar{q}}=g_{\gamma^{*}}^{(0)}\frac{\beta\sqrt{s}}{8\pi}\left(1+\frac{2m^{2}}{s}\right)~{}.$ (17)
@@
$d\Gamma^{(1)}_{f\to q\bar{q}}=\frac{d\Phi_{\bm{\ell}_{1}\bm{\ell}_{2}}}{2\sqrt{s}}\,\bigg{[}\Big{(}{\cal A}_{\rm D}^{(3,f,{\rm R})}(456)\,\widetilde{\Delta}_{45\bar{6}}$
$\quad+{\cal A}_{\rm D}^{(3,f)}(1356)\,\widetilde{\Delta}_{135\bar{6}}\Big{)}+(5\leftrightarrow 2,4\leftrightarrow 3)\bigg{]}~{}.$ (18)
@@
$+2(d-2)\lambda_{2}\left(-\frac{\lambda_{125}\lambda_{1\bar{2}5}}{\lambda_{13\bar{4}}}-\frac{\lambda_{12\bar{5}}\lambda_{1\bar{2}\bar{5}}}{\lambda_{134}}+2\lambda_{1\bar{3}}\right)$
$+\lambda_{1}\left((d-2)\overline{|{\cal M}_{H\to q\bar{q}}^{(0)}|^{2}}-4s\right)\left(\frac{1}{\lambda_{23\bar{4}\bar{5}}}+\frac{1}{\lambda_{2345}}\right)\bigg{\}}$
[…]
@@
$+2(d-2)\lambda_{2}\left(-\frac{\lambda_{125}\lambda_{1\bar{2}5}}{\lambda_{13\bar{4}}}-\frac{\lambda_{12\bar{5}}\lambda_{1\bar{2}\bar{5}}}{\lambda_{134}}+2\lambda_{1\bar{3}}\frac{d-4}{d-2}\right)$
$+\lambda_{1}\left(\left((d-4)\overline{|{\cal M}_{\gamma^{*}\to q\bar{q}}^{(0)}|^{2}}-8s\right)\left(\frac{1}{\lambda_{23\bar{4}\bar{5}}}+\frac{1}{\lambda_{2345}}\right)\right.$
$+\left.\frac{4\lambda_{12\bar{5}}\lambda_{1\bar{2}5}}{\lambda_{23\bar{4}\bar{5}}}+\frac{4\lambda_{125}\lambda_{1\bar{2}\bar{5}}}{\lambda_{2345}}\right)\bigg{\}}+(2\leftrightarrow 3,4\leftrightarrow 5)\bigg{]},$ (20)
@@
$h_{\rm D}\left(456\right)=s(1+\beta^{2})$
×$\times\left[\frac{1}{\lambda_{13\bar{4}}}\left(\frac{1}{\lambda_{23\bar{4}\bar{5}}}+\frac{1}{\lambda_{125}}\right)+\frac{1}{\lambda_{134}\lambda_{2345}}\right]$
$+4\lambda_{2}\bigg{[}\left(\frac{1+\beta^{2}}{2}+\frac{m^{2}}{\lambda_{13\bar{4}}\lambda_{134}}\right)\left(\frac{1}{\lambda_{13\bar{4}}}+\frac{1}{\lambda_{134}}\right)$
$+(d-2)\frac{\lambda_{1\bar{3}}}{s}\bigg{]}~{}.$ (21)
@@
$-\lambda_{134}\lambda_{13\bar{4}}-\frac{m^{2}(\lambda_{125}\lambda_{1\bar{2}5}+\lambda_{134}\lambda_{13\bar{4}})^{2}}{\lambda_{125}\lambda_{1\bar{2}5}\lambda_{134}\lambda_{13\bar{4}}}\bigg{)}\overline{|{\cal M}_{f\to q\bar{q}}^{(0)}|^{2}}$
$+(d-2)\left(\frac{\lambda_{134}\lambda_{13\bar{4}}}{\lambda_{125}\lambda_{1\bar{2}5}}+\frac{\lambda_{125}\lambda_{1\bar{2}5}}{\lambda_{134}\lambda_{13\bar{4}}}\right)+c_{\rm D}^{(f)}\bigg{]}~{},$
$c_{\rm D}^{(H)}=2(d-2)~{},\qquad c_{\rm D}^{(\gamma^{*})}=2(d-4)~{}.$ (22)
In this expression the final-state quark is labelled as $3$ and the antiquark as $5$. The exchange of indices in Eq. (18) accounts for the symmetric phase-space residues
@@
Before moving on to the UV renormalization of the phase-space residues in Eq. (19) and Eq. (20) let us comment on their IR and threshold singularities. Both expressions become singular at $\lambda_{13\bar{4}}\to 0$ and this singularity corresponds to a collinear singularity due to the collinear splitting $13\to 4$, where the outgoing quark is labelled as $4$, and the labels $1$ and $3$ denote the internal particles running in the loop. Since the energy of $4$ is limited by energy conservation, the region in which the loop three-momentum $\bm{\ell}_{1}$ generates a collinear singularity is bounded [68]. Another collinear singularity occurs at $\lambda_{12\bar{5}}\to 0$ when the virtual gluon becomes collinear with the antiquark. If the quark and antiquark are massive, these collinear singularities are shadowed by their mass. Nevertheless, these terms integrate to a mass-dependent logarithm, which is potentially large. The soft singularity, corresponding to a soft virtual gluon, occurs at $\lambda_{1}\to 0$.
@@
Both soft and collinear singularities, or quasicollinear large logarithms for massive quarks, cancel locally with the tree-level phase-space residues presented in Eq. (II), which also diverge as, e.g., $1/\lambda_{13\bar{4}}$ where $1$ and $3$ are now external particles, and $4$ is the parent parton of the collinear splitting. The local cancellation of soft, collinear and quasicollinear configurations ensures that the massless limit of a massive implementation is smooth [20] and, unlike the state-of-the-art approach, does not require a new calculation.
@@
There is also the expected unitary threshold singularity occurring at $\lambda_{23\bar{4}\bar{5}}\to 0$ that generates the absorptive (imaginary) part of the loop amplitude. Threshold singularities are integrable but numerically challenging, and often require the implementation of a contour deformation over the integration domain [69, 70, 71] or other alternative methods [72, 73] to numerically stabilize the integrand. They are known to be one of the main limitations of numerical approaches and imply a compromise between accuracy and lengthy computer jobs. Threshold singularities are, however, absent in the sum of all the unitary phase-space residues because the kernel vacuum amplitude is free of this kind of singularities. This is one of the most remarkable properties of LTD causal unitary.
III Local UV renormalisation and mass renormalisation scheme
The phase-space residues with loop remainders are to be renormalised locally by suitable UV counterterms. For this purpose we follow the procedure presented in detail in Refs. [24, 20]. First, we rescale all the on-shell energies involving the loop three-momentum $\bm{\ell}_{1}$, i.e., by the replacement
$\sqrt{\bm{\ell}_{12}^{2}-\imath 0}\to\sqrt{(\rho\,\bm{\ell}_{1}+\bm{\ell}_{2})^{2}+(\rho^{2}-1)\mu_{\rm UV}^{2}-\imath 0}~{},$ (24)

### v2

@@ Abstract
We present the first proof-of-concept application to decay processes at higher perturbative orders of LTD causal unitary, a novel methodology that exploits the causal properties of vacuum amplitudes in the loop-tree duality (LTD) and is directly well-defined in the four physical dimensions of the space-time. The generation of loop- and tree-level contributions to the differential decay rates from a kernel multiloop vacuum amplitude is shown in detail, and explicit expressions are presented for selected processes that are suitable for a lightweight understanding of the method. Specifically, we provide a clear physical interpretation of the local cancellation of soft, collinear and threshold singularities, and of the local renormalisation of ultraviolet singularities. The presentation is illustrated with numerical results that showcase the advantages of the method.
1 Introduction
In a recent paper Ramirez-Uribe:2024rjg , we have proposed a novel representation of differential observables at high-energy colliders that exploits the manifestly causal properties of the loop-tree duality (LTD) at higher perturbative orders and its connections with directed acyclic graph (DAG) configurations in graph theory Aguilera-Verdugo:2020set ; snowmass2020 ; Aguilera-Verdugo:2020kzc ; Ramirez-Uribe:2020hes ; JesusAguilera-Verdugo:2020fsn ; Sborlini:2021owe ; TorresBobadilla:2021ivx ; Bobadilla:2021pvr ; deJesusAguilera-Verdugo:2021mvg ; Ramirez-Uribe:2021ubp ; Benincasa:2021qcb ; Kromin:2022txz ; Clemente:2022nll ; Capatti:2022mly ; Rios-Sanchez:2024xtv ; Ramirez-Uribe:2024wua . This approach, dubbed LTD causal unitary, generalises the method of four-dimensional unsubtraction (FDU) Hernandez-Pinto:2015ysa ; Sborlini:2016gbr ; Sborlini:2016hat ; Prisco:2020kyb ; Driencourt-Mangin:2017gop ; Driencourt-Mangin:2019sfl ; Driencourt-Mangin:2019aix ; Driencourt-Mangin:2019yhu in which, unlike subtraction methods Kunszt:1992tn ; Frixione:1995ms ; Catani:1996jh ; Catani:1996vz ; Catani:2000vq ; Nagy:2003qn ; Weinzierl:2003fx ; Frixione:2004is ; Kilgore:2004ty ; Catani:2007vq ; Catani:2013tia ; GehrmannDeRidder:2005cm ; Gehrmann-DeRidder:2012too ; Daleo:2006xa ; Somogyi:2006da ; Somogyi:2006db ; Somogyi:2008fc ; Somogyi:2009ri ; Bolzoni:2010bt ; Gleisberg:2007md ; Czakon:2010td ; Czakon:2014oma ; Bevilacqua:2013iha ; NigelGlover:2010kwr ; Currie:2013vh ; Bonciani:2015sha ; Gaunt:2015pea ; Boughezal:2010mc ; Boughezal:2011jf ; Boughezal:2015dra ; DelDuca:2016ily ; Magnea:2018hab ; Magnea:2018ebr ; Caola:2019nzf ; Buonocore:2019puv ; Delto:2019asp ; Engel:2019nfw ; Asteriadis:2019dte ; Cieri:2020ikq ; TorresBobadilla:2020ekr ; Bertolotti:2022aih ; Devoto:2023rpv , infrared (IR) and ultraviolet (UV) singularities are locally cancelled directly between loop and tree contributions at the integrand level, while introducing relevant improvements and new features that facilitate an efficient implementation.
The central and novel ingredient of LTD causal unitary is a multiloop vacuum amplitude in the LTD representation that depends on $\Lambda$ loop momenta, $\{\ell_{j}\}_{j=1,\ldots,\Lambda}$. In LTD causal unitary, the contribution from the $k^{\rm{th}}$-order in perturbation theory to the differential decay rate of a particle of mass $m_{a}$ is given by
$d\Gamma^{(k)}_{a}=\frac{d\Lambda}{2m_{a}}\,\sum_{(i_{1}\cdots i_{n}a)\in\Sigma}{\cal A}_{\rm D}^{(\Lambda,{\rm R})}(i_{1}\cdots i_{n}a)\,{\cal O}_{i_{1}\cdots i_{n}}\,\widetilde{\Delta}_{i_{1}\cdots i_{n}\bar{a}}~{},$ (1)
@@ 1 Introduction
where $d\Gamma^{{\rm N}^{k}{\rm LO}}_{a}$ denotes the differential decay rate up to (next-to)k-leading order, and the integration measure
@@
The final states with $n$ particles are then generated from residues, called phase-space residues, of a vacuum amplitude ${\cal A}_{\rm D}^{(\Lambda)}$ on the on-shell energies $q_{i_{s},0}^{(+)}=\sqrt{\mathbf{q}_{i_{s}}^{2}+m_{i_{s}}^{2}-\imath 0}$ of the internal propagators, where $\imath 0$ is the original complex prescription of a Feynman propagator, $\mathbf{q}_{i_{s}}$ the spacial components of the four-momentum $q_{i_{s}}$, and $m_{i_{s}}$ its mass,
${\cal A}_{\rm D}^{(\Lambda,{\rm R})}(i_{1}\cdots i_{n}a)$ $=$ ${\cal A}_{\rm D}^{(\Lambda)}(i_{1}\cdots i_{n}a)-{\cal A}_{{\rm UV}}^{(\Lambda)}(i_{1}\cdots i_{n}a)~{},$ (4)
@@
The function ${\cal O}_{i_{1}\cdots i_{n}}$ defines the observable under consideration by mapping the spatial components of the internal momenta of the dual vacuum amplitude onto the spatial components of the momenta of the external particles. The external on-shell energies are set by the corresponding equation of motion. The default choice ${\cal O}_{i_{1}\cdots i_{n}}=1$ gives the total decay rate after integration.
@@
The LTD causal unitary representation in Eq. (1) involves the sum over all possible phase-space residues of the vacuum amplitude that lead to the desired final states. It is precisely the sum over the phase-space residues that ensures that most of the unique properties of the vacuum amplitude are preserved. Specifically, the vacuum amplitude in LTD is a function of the on-shell energies and is obtained by replacing the Feynman propagators by causal propagators of the form
@@
Since the real part of the on-shell energies is positive by definition, the causal propagators in Eq. (9) cannot become singular. If all the on-shell energies vanish simultaneously for massless particles, the potential soft singularity is screened by the integration measure, which also vanishes in this limit. As a consequence, the vacuum amplitude cannot exhibit soft or collinear singularities, and remarkably threshold singularities are also absent. The absence of singularities in the vacuum amplitude, apart from UV singularities that are accounted for by the local UV counterterm, ensures that the sum over the phase-space residues is also free of soft, collinear and threshold singularities. General proofs of the matching of singularities between phase-space residues have been presented in Ref. Ramirez-Uribe:2024rjg . Therefore, the master expression for the decay rate in Eq. (1) is well defined in the four physical dimensions of the space-time, since all potential nonintegrable singularities locally cancel out between different phase-space residues.
It is important to stress that LTD causal unitary is conceptually different from using forward-scattering amplitudes to generate final states via unitarity cuts Sterman:1978bi ; Sterman:1978bj ; Ellis:1980wv ; Soper:1998ye ; Soper:1999xk ; Capatti:2020xjc  [^1] In Refs. Sterman:1978bi ; Sterman:1978bj , vacuum polarization diagrams refer to those with two external photons.. The forward-scattering method, where initial-state particles are on-shell from the outset, and unitarity cuts involve only final-state particles, is closely related to the well-known optical theorem. In contrast, LTD causal unitary deals with off-shell particles in vacuum amplitudes, i.e. amplitudes without external particles, and unitarity cuts involve both initial and final-state particles. All external particles are set on-shell through a LTD’s unique property, which ensures this condition when the inverse of the corresponding causal propagator vanishes. In the context of forward-scattering amplitudes, the underlying reason advocated for the cancellation of singularities is unitarity Soper:1998ye . In LTD causal unitary the cancellation of singularities occurs because we depart from an amplitude, the vacuum amplitude, which is intrinsically IR safe and threshold free.
@@
The outline of the paper is as follows. In Section 2, we present explicit expressions for the phase-space residues that contribute to the decay processes $H\to q\bar{q}(g)$ and $\gamma^{*}\to q\bar{q}(g)$ at LO and NLO. In Section 3, we construct the corresponding UV counterterms, which implement a local UV renormalisation, and discuss in detail the wave-function renormalisation and the mass renormalisation scheme. In Section 4, we consider a simplified model that has all the main properties of a NNLO calculation and is therefore suitable for presenting LTD causal unitary at NNLO in a lightweight form. In Section 5, we present a numerical implementation of LTD causal unitary at NLO and NNLO based on the phase-space residues presented in Sections 2 to 4, and illustrate the local cancellation of threshold, double- and triple-collinear/quasicollinear singularities. Finally, in Section 6 we draw our conclusions and future prospects.
2 Vacuum amplitudes and decay rates at NLO
@@
$q_{1}=\ell_{12}~{},\quad q_{2}=\ell_{13}~{},\quad q_{3}=\ell_{1}~{},\quad q_{4}=\ell_{2}~{},\quad q_{5}=\ell_{2\bar{3}}~{},\quad q_{6}=\ell_{3}~{}.$ (10)
We use the shorthand notation $\ell_{ij}=\ell_{i}+\ell_{j}$ and $\ell_{i\bar{j}}=\ell_{i}-\ell_{j}$. Nonetheless, it is worth noting that one of the advantages of the LTD representation is that it is independent of the specific labelling of the internal momenta. More precisely, the relationship between the momenta of the internal propagators and the independent loop momenta in Eq. (10) can be modified by a linear shift. Yet, the LTD representation in terms of causal propagators and on-shell energies remains unaffexcted by the choice of the specific loop routing. The momentum $q_{1}$ represents a gluon and is massless, while $q_{2}$ through $q_{5}$ are quarks or antiquarks of mass $m$. The momentum $q_{6}$ is either a Higgs boson or an off-shell photon. To simplify the presentation, we define $x_{i_{1}\cdots i_{n}}=\prod_{s=1}^{n}2q_{i_{s},0}^{(+)}$.
@@
The bar above an index indicates that the corresponding on-shell energy bears a minus sign, e.g. $\lambda_{45\bar{6}}=q_{4,0}^{(+)}+q_{5,0}^{(+)}-q_{6,0}^{(+)}$. For convenience, the average factor over initial-state polarizations is included in ${\cal A}_{\rm D}^{(2,f)}$. In this expression, we have anticipated that $\overline{|{\cal M}_{f\to q\bar{q}}^{(0)}|^{2}}$ is the matrix element squared obtained with the standard approach of squaring the tree-level amplitude for $f\to q\bar{q}$ and averaging over the initial-state polarizations, where the interaction couplings and final-state colour factors are collected in the coefficient $g_{f}^{(0)}$,
@@
$\Gamma^{\rm LO}_{H\to q\bar{q}}=g_{H}^{(0)}\frac{\beta^{3}m_{H}}{8\pi}~{},\qquad\Gamma^{\rm LO}_{\gamma^{*}\to q\bar{q}}=g_{\gamma^{*}}^{(0)}\frac{\beta\sqrt{s}}{8\pi}\left(1+\frac{2m^{2}}{s}\right)~{}.$ (17)
@@
$d\Gamma^{(1)}_{f\to q\bar{q}}$ $=$ $\frac{d\Phi_{\bm{\ell}_{1}\bm{\ell}_{2}}}{2\sqrt{s}}\,\bigg{[}\Big{(}{\cal A}_{\rm D}^{(3,f,{\rm R})}(456)\,\widetilde{\Delta}_{45\bar{6}}+{\cal A}_{\rm D}^{(3,f)}(1356)\,\widetilde{\Delta}_{135\bar{6}}\Big{)}+(5\leftrightarrow 2,4\leftrightarrow 3)\bigg{]}~{}.$ (18)
@@
$\quad+2(d-2)\lambda_{2}\left(-\frac{\lambda_{125}\lambda_{1\bar{2}5}}{\lambda_{13\bar{4}}}-\frac{\lambda_{12\bar{5}}\lambda_{1\bar{2}\bar{5}}}{\lambda_{134}}+2\lambda_{1\bar{3}}\right)$
$\quad+\lambda_{1}\left((d-2)\overline{|{\cal M}_{H\to q\bar{q}}^{(0)}|^{2}}-4s\right)\left(\frac{1}{\lambda_{23\bar{4}\bar{5}}}+\frac{1}{\lambda_{2345}}\right)\bigg{\}}+(2\leftrightarrow 3,4\leftrightarrow 5)\bigg{]}~{},$ (19)
@@
$\quad+2(d-2)\lambda_{2}\left(-\frac{\lambda_{125}\lambda_{1\bar{2}5}}{\lambda_{13\bar{4}}}-\frac{\lambda_{12\bar{5}}\lambda_{1\bar{2}\bar{5}}}{\lambda_{134}}+2\lambda_{1\bar{3}}\frac{d-4}{d-2}\right)$
$\quad+\lambda_{1}\left(\left((d-4)\overline{|{\cal M}_{\gamma^{*}\to q\bar{q}}^{(0)}|^{2}}-8s\right)\left(\frac{1}{\lambda_{23\bar{4}\bar{5}}}+\frac{1}{\lambda_{2345}}\right)\right.$
$\quad+\left.\frac{4\lambda_{12\bar{5}}\lambda_{1\bar{2}5}}{\lambda_{23\bar{4}\bar{5}}}+\frac{4\lambda_{125}\lambda_{1\bar{2}\bar{5}}}{\lambda_{2345}}\right)\bigg{\}}+(2\leftrightarrow 3,4\leftrightarrow 5)\bigg{]},$ (20)
@@
$h_{\rm D}\left(456\right)=s(1+\beta^{2})\left[\frac{1}{\lambda_{13\bar{4}}}\left(\frac{1}{\lambda_{23\bar{4}\bar{5}}}+\frac{1}{\lambda_{125}}\right)+\frac{1}{\lambda_{134}\lambda_{2345}}\right]$
$\quad+4\lambda_{2}\bigg{[}\left(\frac{1+\beta^{2}}{2}+\frac{m^{2}}{\lambda_{13\bar{4}}\lambda_{134}}\right)\left(\frac{1}{\lambda_{13\bar{4}}}+\frac{1}{\lambda_{134}}\right)+(d-2)\frac{\lambda_{1\bar{3}}}{s}\bigg{]}~{}.$ (21)
@@
$\quad-\lambda_{134}\lambda_{13\bar{4}}-\frac{m^{2}(\lambda_{125}\lambda_{1\bar{2}5}+\lambda_{134}\lambda_{13\bar{4}})^{2}}{\lambda_{125}\lambda_{1\bar{2}5}\lambda_{134}\lambda_{13\bar{4}}}\bigg{)}\overline{|{\cal M}_{f\to q\bar{q}}^{(0)}|^{2}}$
$\quad+(d-2)\left(\frac{\lambda_{134}\lambda_{13\bar{4}}}{\lambda_{125}\lambda_{1\bar{2}5}}+\frac{\lambda_{125}\lambda_{1\bar{2}5}}{\lambda_{134}\lambda_{13\bar{4}}}\right)+c_{\rm D}^{(f)}\bigg{]}~{},$ (22)
with $c_{\rm D}^{(H)}=2(d-2)$, and $c_{\rm D}^{(\gamma^{*})}=2(d-4)$. In this expression the final-state quark is labelled as $3$ and the antiquark as $5$. The exchange of indices in Eq. (18) accounts for the symmetric phase-space residues
@@
Before moving on to the UV renormalization of the phase-space residues in Eq. (19) and Eq. (20) let us comment on their IR and threshold singularities. Both expressions become singular at $\lambda_{13\bar{4}}\to 0$ and this singularity corresponds to a collinear singularity due to the collinear splitting $13\to 4$, where the outgoing quark is labelled as $4$, and the labels $1$ and $3$ denote the internal particles running in the loop. Since the energy of $4$ is limited by energy conservation, the region in which the loop three-momentum $\bm{\ell}_{1}$ generates a collinear singularity is bounded Buchta:2014dfa . Another collinear singularity occurs at $\lambda_{12\bar{5}}\to 0$ when the virtual gluon becomes collinear with the antiquark. If the quark and antiquark are massive, these collinear singularities are shadowed by their mass. Nevertheless, these terms integrate to a mass-dependent logarithm, which is potentially large. The soft singularity, corresponding to a soft virtual gluon, occurs at $\lambda_{1}\to 0$.
@@
Both soft and collinear singularities, or quasicollinear large logarithms for massive quarks, cancel locally with the tree-level phase-space residues presented in Eq. (22), which also diverge as, e.g., $1/\lambda_{13\bar{4}}$ where $1$ and $3$ are now external particles, and $4$ is the parent parton of the collinear splitting. The local cancellation of soft, collinear and quasicollinear configurations ensures that the massless limit of a massive implementation is smooth Sborlini:2016hat and, unlike the state-of-the-art approach, does not require a new calculation.
@@
There is also the expected threshold singularity occurring at $\lambda_{23\bar{4}\bar{5}}\to 0$ that generates the absorptive (imaginary) part of the loop amplitude. Threshold singularities are integrable but numerically challenging, and often require the implementation of a contour deformation over the integration domain Becker:2010ng ; Buchta:2015wna ; Capatti:2019edf or other alternative methods Pittau:2021jbs ; Kermanschah:2021wbk to numerically stabilize the integrand. They are known to be one of the main limitations of numerical approaches and imply a compromise between accuracy and lengthy computer jobs. Threshold singularities are, however, absent in the sum of all the unitary phase-space residues because the kernel vacuum amplitude is free of this kind of singularities. This is one of the most remarkable properties of LTD causal unitary.
3 Local UV renormalisation and mass renormalisation scheme
The phase-space residues with loop remainders are to be renormalised locally by suitable UV counterterms. The novelty introduced in Refs. Ramirez-Uribe:2024rjg ; Rios-Sanchez:2024xtv with respect to previous implementations of a local UV renormalisation at one Sborlini:2016hat and two loops Driencourt-Mangin:2019aix (see also Becker:2010ng ; Donati:2013voa ; Cherchiglia:2020iug ; Capatti:2022tit ) is to extract the leading UV behaviour directly in LTD by rescaling the on-shell energies:
$q_{i_{s},0}^{(+)}\to\sqrt{(\rho\,\bm{\ell}_{s}+\bm{\ell}_{r})^{2}+m_{i_{s}}^{2}+(\rho^{2}-1)\mu_{{\rm UV}}^{2}-\imath 0}~{},$ (24)
and then by expanding the phase-space residues ${\cal A}_{\rm D}^{(3,f)}(456)$ in Eq. (19) and Eq. (20) for $\rho\to\infty$. In Eq. (24), we assume $\mathbf{q}_{i_{s}}=\bm{\ell}_{s}+\bm{\ell}_{r}$, with $\bm{\ell}_{s}$ in the UV region, while $|\bm{\ell}_{r}|\ll|\bm{\ell}_{s}|$.